# Boolean-Narayana numbers

### v1

@@ Abstract
We introduce a refinement of Boolean-Catalan numbers and call them Boolean-Narayana numbers. We provide an explicit formula for these numbers, and prove unimodality, log-concavity, and real-roots-only results for their sequences.
@@ 1 Introduction
The number of 0-1-trees on $n$ vertices is denoted by $B(n)$ and is called a Boolean-Catalan number [10]. They count various objects, some of which are discussed in the thesis of Chetak Hossain [10], and some others appear in papers of Colin Defant [7, 8, 9], later to be shown to bijectively correspond to 0-1-trees by the present author [3]. The name ”Boolean-Catalan numbers” can be justified, for example, by saying that binary plane trees are counted by the Catalan numbers, and if a binary plane tree has $m$ vertices with two children, then we associate an element of the Boolean algebra $\{0,1\}^{m}$ to it.
In this paper, we take this concept further. The Narayana numbers are the best-known refinement of the Catalan numbers. Accordingly, they have many equivalent definitions. The one that is closest to our subject defines the Narayana number $N(n,k)$ as the number of binary plane trees on $n$ vertices with k1k-1 right edges. So we define the Boolean-Narayana numbers $BoNa(n,k)$ as the number of 0-1-trees on $n$ vertices that have k1k-1 right edges.
In Section 2, we provide two additional combinatorial interpretations for the Boolean-Narayana numbers. In Section 3, we use the Lagrange Inversion Formula to prove an explicit formula for these numbers. Then in Section 4, we construct an injection to prove that the for fixed $n$, the sequence of the numbers $BoNa(n,k)$, where $k=1,2,\cdots,n$, is unimodal. We follow this by proving in Section 5 that the generating polynomial of that sequence has real roots only, which implies that it is log-concave. In Section 6, we fix $k$, and prove that the infinite sequence $BoNa(n,k)$, where $n=k+1,k+2,\cdots$, is log-concave. That is, we will see that the Boolean-Narayana numbers share the best-known properties of the Narayana numbers. We end by discussing future directions in Section 7, which also explain our original motivation in this family of questions.
@@ 2 A combinatorial interpretation
The number of permutations in $s^{-1}(\operatorname{Av}_{n}(231,312))$ that have exactly k1k-1 descents is $BoNa(n,k)$.
@@
On the other hand, $B(1,1)=1$, the numbers $BoNa(n,k)$ obviously satisfy the analogous recurrence relation
$BoNa(n,k)=BoNa(n-1,k)+BoNa(n-1,k-1)+2\sum_{i=2}^{n-1}\sum_{j\geq 0}BoNa(i-1,j)BoNa(n-i,k-1-j),$ (2)
@@
The number of permutations in $s^{-1}(\operatorname{Av}_{n}(132,312))$ that have exactly k1k-1 descents is $BoNa(n,k)$.
@@
Note that it follows by reflection through a vertical axis that BoNa(n,k)=BoNa(n,nk+1)BoNa(n,k)=BoNa(n,n-k+1). So Propositions 2.1 and 2.2 yield the following.
@@ 3 Enumeration
In this section, we will use the Lagrange Inversion Formula to find an explicit formula for $BoNa(n,k)$. The Lagrange Inversion Formula, in its simplest form, can be found in many graduate-level textbooks. However, in this section, we will use its univariate version on a function that has two variables, and for that reason, we will explain which version we use and how.
@@
As BoNa(n,k)=BoNa(n,nk+1)BoNa(n,k)=BoNa(n,n-k+1), it suffices to compute $BoNa(n,k)$ for the case when $k\leq(n+1)/2$.
@@
$BoNa(n,k)=\frac{1}{n}{n\choose k-1}\sum_{j=0}^{k-1}2^{j}{k-1\choose j}{n-k+1\choose j+1}$.
Proof: Let $t(n,k)=BoNa(n,k+1)$ be the number of 0-1 trees on $n$ vertices with $k$ right edges, and let
$T(z,u)=\sum_{n\geq 1,k\geq 0}t(n,k)z^{n}u^{k}=1+\sum_{n\geq 1,k\geq 1}BoNa(n,k)z^{n}u^{k-1}.$ (5)
@@
$BoNa(n,k)=[z^{n}u^{k-1}]T(z,u)=\frac{1}{n}{n\choose k-1}[y^{n-1}]y^{k-1}\cdot(1+2y)^{k-1}(1+y)^{n-k+1},$
@@
$BoNa(n,k)=[z^{n}u^{k-1}]T(z,u)=\frac{1}{n}{n\choose k-1}[y^{n-k}](1+2y)^{k-1}(1+y)^{n-k+1}.$
@@
$BoNa(n,k)=\frac{1}{n}{n\choose k-1}\sum_{j=0}^{\min(k-1,n-k)}2^{j}{k-1\choose j}{n-k+1\choose j+1}$,
@@
The numbers $BoNa(n,k)$ for $1\leq n\leq 5$, and $k=0,1,\cdots,n-1$ are shown below. Note the interesting fact that $BoNa(n,2)=(n-1)^{2}$.
@@
4 Unimodality for fixed $n$
@@
For all fixed $n$, the sequence $BoNa(n,k)_{1\leq k\leq n}$ is unimodal.
Proof: As for any fixed $n$, the sequence $BoNa(n,k)_{1\leq k\leq n}$ is symmetric, it suffices to show that $BoNa(n,k)\leq B(n,k+1)$ if $k\leq(n-1)/2$. Let $\mathcal{B}oNa(n,k)$ be the set of 0-1 trees with k1k-1 right edges. We will construct an injection $z:\mathcal{B}oNa(n,k)\rightarrow\mathcal{B}oNa(n,k+1)$. Note that our injection $z$ is a slightly modified version that the present author used to prove a unimodality result on $t$-stack sortable permutations in [4, 5]. A perhaps more accessible source is Section 8.2.3 of [2].
@@
Figure 3 shows an example of the injection $z:\mathcal{B}(9,4)\rightarrow\mathcal{B}(9,5)$. The large numbers denote the label of vertices with two children, whereas the small labels show the place of a vertex in the total order we defined above. In this example, the smallest index $i$ so that $\ell(T_{i})-r(T_{i})=1$ is $i=3$.
Figure 3: The action of $z$ on a tree in $\mathcal{B}(9,4)$.
@@
For all $k\leq(n-1)/2$, the function $z$ described above is an injection from $\mathcal{B}(n,k)$ into $\mathcal{B}(n,k+1)$.
Proof: It is clear that $z$ maps into $\mathcal{B}(n,k+1)$ as $z$ consists of the application of $f$ to a subgraph $T_{i}$ in which left edges outnumber right edges by one. We know that $f$ turns the number of left edges into the number of right edges and vice versa, so $z$ indeed increases the number of right edges by 1.
@@
Now let $U\in\mathcal{B}(n,k+1)$. If $U$ does not have a subforest that consists of the first few nodes of $U$ in the total order of all nodes and in which the right edges outnumber the left edges by one, then by the previous paragraph, $U$ has no preimage under $z$. Otherwise, the unique preimage $z^{-1}(U)$ can be found by finding the smallest such subforest, applying $f$ to it, and leaving the rest of $U$ unchanged. $\Diamond$
Lemma 4.3 shows that $BoNa(n,k)\leq BoNa(n,k+1)$ if $k\leq(n-1)/2$. This implies that the sequence $\{BoNa(n,k)\}_{1\leq k\leq n}$ weakly increases to its midpoint. As the sequence is symmetric, it follows that it weakly decreases after that. Therefore, the sequence is unimodal as claimed. $\Diamond$
@@ 5 The real zeros property
Let $BoNa_{n}(u)=\sum_{k=1}^{n}BoNa(n,k)u^{k}$. We call these polynomials the Boolean-Narayana polynomials.
@@
For all positive integers $n$, the Boolean-Narayana polynomial $BoNa_{n}(u)$ has real roots only.
Proof: Note that $BoNa_{n}(u)=[z^{n}](uT(z,u))$. Note that for a fixed $u$, (8) implies, using (6), that
@@
$\sum_{k\geq 1}^{n}BoNa(n,k)u^{k-1}=[z^{n}]T(z,u)=\frac{1}{n}[y^{n-1}]\left(1+(1+u)y+2uy^{2}\right)^{n}.$
@@
$BoNa_{n}(x)=\frac{u}{n}[y^{n-1}]\left(1+(1+u)y+2uy^{2}\right)^{n}.$ (9)
@@
The numbers ${n\choose j}{n\choose j+1}$ are constant multiples of the Narayana numbers $N(n,j)$. Therefore, we define the Narayana polynomials as
$N_{n}(q)=\frac{1}{n}\sum_{j=0}^{n-1}{n\choose j}{n\choose j+1}q^{j}.$ (11)
@@
$BoNa_{n}(u)=u(-1)^{n-1}r_{2}^{n-1}N_{n}(q(u)).$
It was proved by Francesco Brenti (see Theorem 5.3.1 in [6]) that all n1n-1 roots of $N_{n}(u)$ are real and negative. It is also known that the roots are simple (see Theorem 3.2 in [11]). The last displayed equation shows that if $q(u)$ is a root of $N_{n}$, then $u$ is a root of $BoNa_{n}$. As both polynomials have degree $n$ and have a root equal to 0, this will complete the proof of our theorem if we can show that the function $q(u)$ is bijective from the set $(-\infty,0)$ to itself. In order to do that we compute
$q(u)=\frac{r_{1}(u)}{r_{2}(u)}=\frac{1+u-\Delta}{1+u+\Delta},$
where $\Delta=\sqrt{1-6u+u^{2}}$. Then
$q^{\prime}(u)=\frac{-8(u-1)}{(u+1+\Delta)^{2}\Delta}.$
Now note that if $u<0$, then the 8(u1)=8(1u)>0-8(u-1)=8(1-u)>0, and $(u+1+\Delta)^{2}>0$, so $q^{\prime}(u)>0$, and so $q(u)$ is an increasing function on the interval $(-\infty,0)$. Finally, $q(u)$ takes all negative real values, since it is a continuous function on the set of negative reals and $\lim_{u\rightarrow-\infty}q(u)=$-$\infty$ and $\lim_{u\rightarrow 0}q(u)=0$. $\Diamond$
@@
For all fixed $n$, the sequence $BoNa(n,1),BoNa(n,2),\cdots,BoNa(n,n)$ is log-concave.

### v3

@@ Abstract
We introduce a refinement of Boolean-Catalan numbers and call them Boolean-Narayana numbers. We provide an explicit formula for these numbers, and prove unimodality, log-concavity, and real-roots-only results for their sequences. We also prove a three-term recurrence relation for their generating polynomials.
@@ 1 Introduction
The number of 0-1-trees on $n$ vertices is denoted by $B(n)$ and is called a Boolean-Catalan number [11]. The Boolean-Catalan numbers count various objects, some of which are discussed in the thesis of Chetak Hossain [11], and some others appear in papers of Colin Defant [7, 8, 9], later to be shown to bijectively correspond to 0-1-trees by the present author [3]. The name ”Boolean-Catalan numbers” can be justified, for example, by saying that binary plane trees are counted by the Catalan numbers, and if a binary plane tree has $m$ vertices with two children, then we associate an element of the Boolean algebra $\{0,1\}^{m}$ to it.
In this paper, we take this concept further. The Narayana numbers are the best-known refinement of the Catalan numbers. Accordingly, they have many equivalent definitions. The one that is closest to our subject defines the Narayana number $N(n,k)$ as the number of binary plane trees on $n$ vertices with k1k-1 right edges. So we define the Boolean-Narayana numbers $\mathrm{BoNa}(n,k)$ as the number of 0-1-trees on $n$ vertices that have k1k-1 right edges.
In Section 2, we provide two additional combinatorial interpretations for the Boolean-Narayana numbers. In Section 3, we use the Lagrange Inversion Formula to prove an explicit formula for these numbers. Then in Section 4, we construct an injection to prove that the for fixed $n$, the sequence of the numbers $\mathrm{BoNa}(n,k)$, where $k=1,2,\cdots,n$, is unimodal. We follow this by proving in Section 5 that the generating polynomial $\mathrm{BoNa}_{n}(u)$ of that sequence has real roots only, which implies the sequence of its coefficient is log-concave. We also prove that the polynomials $\mathrm{BoNa}_{n}(u)$ have weakly interlacing roots. In Section 6, we prove a three-term recurrence relation satisfied by the polynomials $\mathrm{BoNa}_{n}(u)$. Then in Section 7 we use this recurrence relation to provide an alternative proof of that interlacing property of these polynomials. In Section 8, we fix $k$, and prove that the infinite sequence $\mathrm{BoNa}(n,k)$, where $n=k+1,k+2,\cdots$, is log-concave.
That is, we will see that the Boolean-Narayana numbers share the best-known properties of the Narayana numbers. This is interesting, since the explicit formula we obtain for these numbers is not as compact as the well-known formula $N(n,k)=\frac{1}{n}{n\choose k}{n\choose k+1}$ for the latter. We end by discussing future directions in Section 9, where we also explain our original motivation for this family of questions.
@@ 2 A combinatorial interpretation
The number of permutations in $s^{-1}(\operatorname{Av}_{n}(231,312))$ that have exactly k1k-1 descents is $\mathrm{BoNa}(n,k)$.
@@
On the other hand, $\mathrm{BoNa}(1,1)=1$, and the numbers $\mathrm{BoNa}(n,k)$ obviously satisfy the analogous recurrence relation
$\mathrm{BoNa}(n,k)=\mathrm{BoNa}(n-1,k)+\mathrm{BoNa}(n-1,k-1)+2\sum_{i=2}^{n-1}\sum_{j\geq 0}\mathrm{BoNa}(i-1,j)\mathrm{BoNa}(n-i,k-1-j),$ (2)
@@
The number of permutations in $s^{-1}(\operatorname{Av}_{n}(132,312))$ that have exactly k1k-1 descents is $\mathrm{BoNa}(n,k)$.
@@
Note that it follows by reflection through a vertical axis that $\mathrm{BoNa}(n,k)=\mathrm{BoNa}(n,n-k+1)$. So Propositions 2.1 and 2.2 yield the following.
@@ 3 Enumeration
In this section, we will use the Lagrange Inversion Formula to find an explicit formula for $\mathrm{BoNa}(n,k)$. The Lagrange Inversion Formula, in its simplest form, can be found in many graduate-level textbooks. However, in this section, we will use its univariate version on a function that has two variables, and for that reason, we will explain which version we use and how.
@@
As $\mathrm{BoNa}(n,k)=\mathrm{BoNa}(n,n-k+1)$, it suffices to compute $\mathrm{BoNa}(n,k)$ for the case when $k\leq(n+1)/2$.
@@
$\mathrm{BoNa}(n,k)=\frac{1}{n}{n\choose k-1}\sum_{j=0}^{k-1}2^{j}{k-1\choose j}{n-k+1\choose j+1}$.
Proof: Let $t(n,k)=\mathrm{BoNa}(n,k+1)$ be the number of 0-1 trees on $n$ vertices with $k$ right edges, and let
$T(z,u)=\sum_{n\geq 1,k\geq 0}t(n,k)z^{n}u^{k}=1+\sum_{n\geq 1,k\geq 1}\mathrm{BoNa}(n,k)z^{n}u^{k-1}.$ (5)
@@
$\mathrm{BoNa}(n,k)=[z^{n}u^{k-1}]T(z,u)=\frac{1}{n}{n\choose k-1}[y^{n-1}]y^{k-1}\cdot(1+2y)^{k-1}(1+y)^{n-k+1},$
@@
$\mathrm{BoNa}(n,k)=[z^{n}u^{k-1}]T(z,u)=\frac{1}{n}{n\choose k-1}[y^{n-k}](1+2y)^{k-1}(1+y)^{n-k+1}.$
@@
$\mathrm{BoNa}(n,k)=\frac{1}{n}{n\choose k-1}\sum_{j=0}^{\min(k-1,n-k)}2^{j}{k-1\choose j}{n-k+1\choose j+1}$,
@@
The numbers $\mathrm{BoNa}(n,k)$ for $1\leq n\leq 5$, and $k=0,1,\cdots,n-1$ are shown below. Note the interesting fact that $\mathrm{BoNa}(n,2)=(n-1)^{2}$.
@@
4 A Combinatorial Proof for Unimodality for fixed $n$
@@
For all fixed $n$, the sequence $\mathrm{BoNa}(n,k)_{1\leq k\leq n}$ is unimodal.
Proof: As for any fixed $n$, the sequence $\mathrm{BoNa}(n,k)_{1\leq k\leq n}$ is symmetric, it suffices to show that $\mathrm{BoNa}(n,k)\leq\mathrm{BoNa}(n,k+1)$ if $k\leq(n-1)/2$. Let $\mathcal{B}oNa(n,k)$ be the set of 0-1 trees with k1k-1 right edges. We will construct an injection $z:\mathcal{B}oNa(n,k)\rightarrow\mathcal{B}oNa(n,k+1)$. Note that our injection $z$ is a slightly modified version that the present author used to prove a unimodality result on $t$-stack sortable permutations in [4, 5]. A perhaps more accessible source is Section 8.2.3 of [2].
@@
Figure 3 shows an example of the injection $z:\mathcal{B}oNa(9,4)\rightarrow\mathcal{B}oNa(9,5)$. The large numbers denote the label of vertices with two children, whereas the small labels show the place of a vertex in the total order we defined above. In this example, the smallest index $i$ so that $\ell(T_{i})-r(T_{i})=1$ is $i=3$.
Figure 3: The action of $z$ on a tree in $\mathcal{B}oNa(9,4)$.
@@
For all $k\leq(n-1)/2$, the function $z$ described above is an injection from $\mathcal{B}oNa(n,k)$ into $\mathcal{B}oNa(n,k+1)$.
Proof: It is clear that $z$ maps into $\mathcal{B}oNa(n,k+1)$ as $z$ consists of the application of $f$ to a subgraph $T_{i}$ in which left edges outnumber right edges by one. We know that $f$ turns the number of left edges into the number of right edges and vice versa, so $z$ indeed increases the number of right edges by 1.
@@
Now let $U\in\mathcal{B}oNa(n,k+1)$. If $U$ does not have a subforest that consists of the first few nodes of $U$ in the total order of all nodes and in which the right edges outnumber the left edges by one, then by the previous paragraph, $U$ has no preimage under $z$. Otherwise, the unique preimage $z^{-1}(U)$ can be found by finding the smallest such subforest, applying $f$ to it, and leaving the rest of $U$ unchanged. $\Diamond$
Lemma 4.3 shows that $\mathrm{BoNa}(n,k)\leq\mathrm{BoNa}(n,k+1)$ if $k\leq(n-1)/2$. This implies that the sequence $\{\mathrm{BoNa}(n,k)\}_{1\leq k\leq n}$ weakly increases to its midpoint. As the sequence is symmetric, it follows that it weakly decreases after that. Therefore, the sequence is unimodal as claimed. $\Diamond$
@@ 5 The real zeros property
Let $\mathrm{BoNa}_{n}(u)=\sum_{k=1}^{n}\mathrm{BoNa}(n,k)u^{k}$. We call these polynomials the Boolean-Narayana polynomials.
@@
For all positive integers $n$, the Boolean-Narayana polynomial $\mathrm{BoNa}_{n}(u)$ has real roots only.
Proof: Note that $\mathrm{BoNa}_{n}(u)=[z^{n}](uT(z,u))$. Note that for a fixed $u$, (8) implies, using (6), that
@@
$\sum_{k\geq 1}^{n}\mathrm{BoNa}(n,k)u^{k-1}=[z^{n}]T(z,u)=\frac{1}{n}[y^{n-1}]\left(1+(1+u)y+2uy^{2}\right)^{n}.$
@@
$\mathrm{BoNa}_{n}(u)=\frac{u}{n}[y^{n-1}]\left(1+(1+u)y+2uy^{2}\right)^{n}.$ (9)
@@
Crucially, and that is why (11) is important, the numbers ${n\choose j}{n\choose j+1}$ are constant multiples of the Narayana numbers $N(n,j)$. Therefore, we define the Narayana polynomials as
$N_{n}(q)=\sum_{j=1}^{n}{n\choose j}{n\choose j-1}q^{j}.$ (12)
@@
$\mathrm{BoNa}_{n}(u)=u(-1)^{n-1}r_{2}^{n-1}N_{n}(q(u)).$ (13)
Equation (13) shows the close connection between the polynomials $\mathrm{BoNa}_{n}(u)$ and $N_{n}(q)$. It is the basis for our results in the rest of this section and the next.
It is known [10] that all n1n-1 nonzero roots of $N_{n}(u)$ are real and simple. These roots are clearly all negative. Other sources for these results include [6] (Theorem 5.3.1) and [13] (Theorem 3.2.) The last displayed equation shows that if $q(u)$ is a root of $N_{n}$, then $u$ is a root of $\mathrm{BoNa}_{n}$. As both polynomials have degree $n$ and have a root equal to 0, this will complete the proof of our theorem if we can show that the function $q(u)$ is bijective from the set $(-\infty,0)$ to itself. Set $\Delta=1-6u+u^{2}$. As $\Delta>0$ if $u<0$, the square root of $\Delta$ is defined as a real number, and we set $s=\sqrt{\Delta}$ for $u<0$.
Now note that
$q(u)=\frac{r_{1}(u)}{r_{2}(u)}=\frac{1+u-s}{1+u+s}.$
Therefore,
$q^{\prime}(u)=\frac{-8(u-1)}{(u+1+s^{2})s}.$
Now note that if $u<0$, then 8(u1)=8(1u)>0-8(u-1)=8(1-u)>0, and $(u+1+s)^{2}>0$, so $q^{\prime}(u)>0$, and so $q(u)$ is an increasing function on the interval $(-\infty,0)$. Finally, $q(u)$ takes all negative real values, since it is a continuous function on the set of negative reals and $\lim_{u\rightarrow-\infty}q(u)=$-$\infty$ and $\lim_{u\rightarrow 0}q(u)=0$. $\Diamond$
@@
For all fixed $n$, the sequence $\mathrm{BoNa}(n,1),\mathrm{BoNa}(n,2),\cdots,\mathrm{BoNa}(n,n)$ is log-concave.
@@
We return to the function $q(u)$ and collect our observations about the function $q(u)$ in the following corollary for future use.
Corollary 5.3
If $u<0$, then $q(u)$ is a negative real number. In fact, the function
$q:(-\infty,0)\to(-\infty,0)$
is continuous and strictly increasing. As $q$ takes all real negative values, it is a bijection.
We are now ready to prove our next theorem.
Theorem 5.4
The polynomials $\{\mathrm{BoNa}_{n}(u)\}_{n\geq 2}$ have weakly interlacing roots. That is, each such polynomial has 0 as a root, and between any two consecutive roots $v_{i}$ and $v_{i+1}$ of $\mathrm{BoNa}_{n}(u)$, there is exactly one root $w_{i}$ of $\mathrm{BoNa}_{n-1}(u)$.
Note that another usual way of expressing the fact the interlacing property of the polynomials $\mathrm{BoNa}_{n}(u)$ is by saying that these polynomials form a (weak) Sturm sequence.
Proof: It is well-known [10] that the Narayana polynomials $\{N_{n}(q)\}_{n\geq 2}$ have weakly interlacing real zeros. Let $\alpha_{i}$ and $\alpha_{i+1}$ be two consecutive negative roots of $N_{n}(q)$, and let $\beta_{i}$ be the unique root of $N_{n-1}(q)$ that is located between them, that is, for which $\alpha_{i}<\beta_{i}<\alpha_{i+1}$ holds.
Corollary 5.3 shows that $q$ is an order-preserving bijection from $(-\infty,0)$ to $(-\infty,0)$; therefore, so is $q^{-1}$. So $\alpha_{i}<\beta_{i}<\alpha_{i+1}$ implies that $v_{i}=q^{-1}(\alpha_{i})<w_{i}=q^{-1}(\beta_{i})<v_{i+1}=q^{-1}(\alpha_{i+1})$. Furthermore, (13) shows that the roots of $\mathrm{BoNa}_{n}(u)$ are precisely the preimages of the roots of $N_{n}$ under $q$, and they are in the same order as $q^{-1}$ is order-preserving. So root $w_{i}$ of $\mathrm{BoNa}_{n-1}(u)$ is between two consecutive roots ( $v_{i}$ and $v_{i+1}$) of $\mathrm{BoNa}_{n}(u)$, and our proof is complete. $\Diamond$
6 A three-term recurrence
In this section, we use a known three-term recurrence for the Narayana polynomials and the relation (13) between the Narayana polynomials and the Boolean-Narayana poylynomials to prove a three-term recurrence for the latter.
Recall that we defined $r_{1}(u)$ and $r_{2}(u)$ right after (10). An equivalent definition is that they are the two roots of
$t^{2}+(1+u)t+2u=0.$ (14)
Also recall that this means that $r_{1}(u)+r_{2}(u)=-(1+u)$, and $r_{1}(u)r_{2}(u)=2u$, and keep the notation $q(u)=\frac{r_{1}(u)}{r_{2}(u)}$.
We will use the standard three-term recurrence for the Narayana polynomials [10], which states that
$(n+1)N_{n}(x)=(2n-1)(1+x)N_{n-1}(x)-(n-2)(x-1)^{2}N_{n-2}(x)$ (15)
if $n\geq 3$.
Substitute $x=q(u)$ in (15) and multiply both sides by $u(-1)^{n-1}r_{2}(u)^{n-1}$. Using (13) that describes the relation between the polynomials $\mathrm{BoNa}_{n}$ and $N_{n}$, we obtain
$(n+1)\mathrm{BoNa}_{n}(u)$ $=(2n-1)\,u(-1)^{n-1}r_{2}^{n-1}\bigl(1+q\bigr)\,N_{n-1}(q)-(n-2)\,u(-1)^{n-1}r_{2}^{n-1}\bigl(q-1\bigr)^{2}\,N_{n-2}(q)$
$=-(2n-1)\,r_{2}(u)\bigl(1+q(u)\bigr)\,\mathrm{BoNa}_{n-1}(u)-(n-2)\,r_{2}(u)^{2}\bigl(q(u)-1\bigr)^{2}\,\mathrm{BoNa}_{n-2}(u).$ (16)
The following identities enable us to eliminate $r_{1}$, $r_{2}$, and $q$ from the obtained recurrence. First,
$r_{2}(1+q)=-r_{2}\left(1+\frac{r_{1}}{r_{2}}\right)=-(r_{1}+r_{2})=1+u.$
Second,
$r_{2}^{2}(q-1)^{2}=r_{2}^{2}\left(\frac{r_{1}-r_{2}}{r_{2}}\right)^{2}=(r_{1}-r_{2})^{2}.$
Using (14),
$(r_{1}-r_{2})^{2}=(r_{1}+r_{2})^{2}-4r_{1}r_{2}=(1+u)^{2}-8u=u^{2}-6u+1.$
Substituting these into (16) yields the three-term recurrence that was our goal to prove in this section.
Theorem 6.1
For all $n\geq 3$, the recurrence relation
$(n+1)\,\mathrm{BoNa}_{n}(u)=(2n-1)(1+u)\,\mathrm{BoNa}_{n-1}(u)-(n-2)(u^{2}-6u+1)\,\mathrm{BoNa}_{n-2}(u),$ (17)
holds.
With initial values
$\mathrm{BoNa}_{1}(u)=u,\qquad\mathrm{BoNa}_{2}(u)=u+u^{2},$
the recurrence (17) determines all polynomials $\mathrm{BoNa}_{n}(u)$.
7 Proving the interlacing property from the three-term recurrence
Theorem 6.1 and a result of Liu and Wang [12] yield an alternative proof for Theorem 5.4 that states that the polynomials $\mathrm{BoNa}_{n}(u)$ have weakly interlacing roots. The mentioned result from [12] is more general than what we need here, and we only cite the case that we will use.
Lemma 7.1
(Corollary 2.4 in [12]) Let $\{P_{n}(x)\}_{n}$ be a sequence of polynomials with positive coefficients that satisfies the recurrence relation
$P_{n}(x)=a_{n}(x)P_{n-1}(x)+b_{n}(x)P_{n-1}^{\prime}(x)+c(x)P_{n-2}(x),$ (18)
where $a_{n}$, $b_{n}$ and $c_{n}$ are polynomials with real coefficients so that $\mathrm{deg}P_{n}=\mathrm{deg}P_{n-1}+1$. If $b_{n}(x)\leq 0$ and $c_{n}(x)\leq 0$ whenever $x\leq 0$, and at least one of $b_{n}(x)<0$ and $c_{n}(x)<0$ holds for all $n$, then the roots of $P_{n}(x)$ and $P_{n-1}(x)$ are weakly interlacing.
Here is the alternative proof of Theorem 5.4 that we promised.
Proof: Let $n\geq 3$. Divide (17) by $(n+1)$ to obtain
$\mathrm{BoNa}_{n}(u)=a_{n}(u)\,\mathrm{BoNa}_{n-1}(u)+b_{n}(u)\,\mathrm{BoNa}^{\prime}_{n-1}(u)+c_{n}(u)\,\mathrm{BoNa}_{n-2}(u),$
where
$a_{n}(u)=\frac{2n-1}{n+1}(1+u),\qquad b_{n}(u)=0,\qquad c_{n}(u)=-\frac{n-2}{n+1}(u^{2}-6u+1).$
Since $b_{n}(u)\equiv 0$, we have $b_{n}(u)\leq 0$ for all $u\leq 0$. Moreover, for $u\leq 0$ one has $u^{2}-6u+1=u^{2}+(-6u)+1\geq 1$, hence for $n\geq 3$, we have
$c_{n}(u)=-\frac{n-2}{n+1}(u^{2}-6u+1)<0\qquad\text{for all }u\leq 0.$
By construction, each $\mathrm{BoNa}_{n}(u)$ has nonnegative coefficients, and $\deg\mathrm{BoNa}_{n}=n$. Therefore the hypotheses of Lemma 7.1 are satisfied, and it follows that $\{\mathrm{BoNa}_{n}(u)\}$ is a Sturm sequence: each $\mathrm{BoNa}_{n}$ is real-rooted with simple zeros and the nonzero roots of $\mathrm{BoNa}_{n-1}$ strictly interlace the nonzero roots of $\mathrm{BoNa}_{n}$ for all $n\geq 2$. $\Diamond$